OpenStudy (anonymous):

Why does the derivative of x^3/2 + 1 = 3/2x^1/2?

OpenStudy (anonymous):

\[ x^{3/2}+1 = \frac{ 3 }{ 2 }x ^{1/2}\]

OpenStudy (anonymous):

I understand by the power rule that i pull the 3/2s out front and then subtract from the exponent and get the 3/2x^1/2, but where does the 1 go?

OpenStudy (anonymous):

the derivative of one is zero

OpenStudy (anonymous):

always?

OpenStudy (anonymous):

adding one to the end raises the function up by one unit,but it does not change its shape the slopes are still the same

OpenStudy (anonymous):

the derivative of a constant is zero, just like the slope of a horizontal line is zero

OpenStudy (anonymous):

well my book shows the derivative of 2+x^-1 = -x^-2 Where did the 2 go?

OpenStudy (anonymous):

two is a constant

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (aravindg):

"derivative of constant =0"

OpenStudy (anonymous):

so if i'm taking the dervative of any constant number than its always 0?

OpenStudy (aravindg):

see above ^

OpenStudy (anonymous):

thanks! would have been nice for my teacher to tell me that!

OpenStudy (anonymous):

think of the graph of \(y=x^2\) and the graph of \(y=x^2+2\) the shapes are identical, the second is just two units above the first

OpenStudy (anonymous):

yes, the derivative of any constant is zero, so the derivative of \(x^2\) and \(x^2+3\) and \(x^2-1\) are all \(2x\)

OpenStudy (anonymous):

Thanks!

OpenStudy (aravindg):

think of it like this dy/dx measures change of y with respect to x..if y= a constant then there is no change at all . so derivative=0